Calculate the mean and the variance of the discrete random variable X which can take only values 1, 2, and 3, given that P(1) = 10/33, P(2) = 1/3, and P(3) = 12/33.




P(X) X-P(X) X²P(X)


























Solve for the mean and the variance of the discrete random variable X which can take only the values 2, 4, 5 and 9, given that P(2) = 9/20, P(4) = 1/20, P(5) = 1/5 and P(9) = 3/10.



X P(X) X∙P(X) X2∙P(X)




























1
Expert's answer
2022-03-15T12:11:27-0400

a)

Mean of X.

"\mu"x="\sum" X.P(X)

=1(10/33)+2(1/3)+3(12/33)

=2.06

Variance of X.

Var(X) ="\sum" (X-"\mu"x)2.P(X)

=(1-2.06)2(10/33)+(2-2.06)2(1/3)+(3-2.06)2(12/33)

=0.663


b)


Mean of X.


"\mu"x=∑ X.P(X)


=2(9/20)+4(1/20)+5(1/5)+9(3/10)


=4.8


Variance of X.


Var(X) =∑ (X-"\mu")2.P(X)


=(2-4.8)2(9/20)+(4-4.8)2(1/20)+(5-4.8)2(3/10)


=8.86